Jennifer is 13 years older than her cousin, Jane, and three times as old as her cousin Freddy. Jane is twice as old as her little brother, Jack. If Jack is one year old, then how old is Jennifer?

Answer by lwsshak3(11628)

Jennifer is 13 years older than her cousin, Jane, and three times as old as her cousin Freddy. Jane is twice as old as her little brother, Jack. If Jack is one year old, then how old is Jennifer?
***
Jack's age=1
Jane's age=2*1=2
Jennifer's age=13+2=15